Hello, I am requesting help with #4.... I'm a little confused on what the question is asking. Thank you in advance.

Communications Unlimited provides support services to its clients.
This company expects to earn an annual return on the assets invested at a rate of: 16%
The company has the following amount invested in the business: $6,000,000
The annual budgeted costs for next year are:
Variable costs Fixed costs
Support Services $400,000 1,900,000
The annual budgeted hours for next year are:
Consulting services                        50,000 hours
Required: You must use cell references for all calculations.
1.  Determine the Target profit for the company.
$                                                                            960,000.00
2.  Determine the markup on variable costs in percentage terms.
$                                                                                       1.40
3.  Determine the  variable cost per hour.
$                                                                                       8.00
4.  Determine the revenue per hour that will be charged if variable costs is the basis for markup.
5.  Determine the markup on total costs in percentage terms. Variable + Fixed
42% $2,300,000
Answer #1
Please give positive ratings so I can keep answering. Thanks!
Communications Unlimited
Answer a
Amount invested in the business    6,000,000.00 A
Desired return 16% B
Target Profit       960,000.00 C=A*B
Total Mark up
Desired return       960,000.00 See C
Total Fixed costs    1,900,000.00 D
Total Mark up 2,860,000.00 E=C+D
Answer 1
Total Variable costs       400,000.00 F
Total Mark up    2,860,000.00 See E
Markup on Total Variable costs 715.00% G=E/F
Answer 2
Total Variable costs       400,000.00 See F
Annual budgeted hours         50,000.00 H
Variable costs per hour                   8.00 I=F/H
Answer 3
Total Variable costs       400,000.00 See F
Total Mark up    2,860,000.00 See E
Total Revenue 3,260,000.00 J=F+E
Annual budgeted hours         50,000.00 See H
Revenue per hour                 65.20 K=J/H
Answer 4
Total Variable costs       400,000.00 See F
Total Fixed costs    1,900,000.00 See D
Total costs 2,300,000.00 L=F+D
Desired return       960,000.00 See C
Markup on Total costs 42.00% M=C/L